OpenStudy (anonymous):

Celeste wants to have her hair cut and permed and also go to lunch. She knows she will need $59. The perm costs twice as much as her haircut and she needs $5 for lunch. How much does the perm cost? @beccaboo333 @AllTehMaffs

OpenStudy (beccaboo333):

Whaaa?

OpenStudy (anonymous):

Lol that's what I said ._.

OpenStudy (beccaboo333):

Sorry i'm not great with math.

OpenStudy (anonymous):

so if you use x as the variable, can you define the haircut and the perm in terms of x?

OpenStudy (anonymous):

Thanks @beccaboo333 however I'll wait until AllTehMaffs explains it to me as-for I feel he'll be explaining xD.

OpenStudy (beccaboo333):

Cost of Haircut = x Cost of Perm = 2x Cost of Lunch = $5 x + 2x + 5 = 59 3x = 59 - 5 3x = 54 x = 54 / 3 x = 18 Cost of Perm = $36

OpenStudy (beccaboo333):

Am I right? @AllTehMaffs

OpenStudy (anonymous):

Thanks. And lol guess so.

OpenStudy (anonymous):

yes, that's the correct answer

OpenStudy (anonymous):

Yeah, you are.

OpenStudy (beccaboo333):

I feel good now lol, considering this is the one subject i really am not good in